OpenStudy (anonymous):

the quantity 4+radical 8 times the quantity 4 minus 4 radical 8.

OpenStudy (anonymous):

\[{{(4+\sqrt{8})} \times {(4-{4}\sqrt{8})}}\]

OpenStudy (anonymous):

is this your question?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

plz see the second term once

OpenStudy (anonymous):

take to second 4 away ,just leave one 4

OpenStudy (anonymous):

\[{4^2 - \sqrt{8^2}}= 16 - 8 = 8\]

OpenStudy (anonymous):

(4 + sqrt8)*(4-sqrt8) = 4^2 -(sqrt8)^2 = 16 - 8 = 8

OpenStudy (anonymous):

thanks :)